OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y= e^x/x, (1,e). I know that these are simpler than they look. I'm having difficulty keeping the formulas straight.

OpenStudy (anonymous):

So first differentiate the equation of the curve. \[y = (e^x)/x\] \[y' = (e^x)/x - (e^x)/x^2\]

OpenStudy (anonymous):

To solve for the slope, substitute 1 in place of x, since it is the x-coordinate of the given point. \[m = (e^1)/1 - (e^1)/1^2\] \[m = e-e\] \[m = 0\]

OpenStudy (anonymous):

Use the point-slope formula to find the equation of the tangent line. y - y1=m(x - x1) y - e = 0(x - 1) y = e

OpenStudy (anonymous):

Ugh, I knew it was simple! Thanks for the help! I feel silly now!

OpenStudy (anonymous):

np